Please help me with this word problem!
A rectangle has a perimeter of 6 centimeters and an area of 2 square centimeters. Find its dimensions.

The answer is 2 centimeters by 1 centimeter but I don't know how to get that.

Answer by sudhanshu_kmr(1152)

Let length = x, width = y
perimeter = 2(x+y) = 6
and area = x*y = 2

as, 2(x+y) = 6
=> x+y = 3
=> x = 3-y
after putting the value of x on x*y = 2
we get (3-y)y= 2
=> y^2 -3y +2 =0
=> (y-1)(y-2) = 0 after factorizing
so, y = 1 or 2
if y =1, x =2
y = 2, x =1

so, its dimensions are 2 and 1.